# THE RESTRICTED PICARD FUNCTOR

### v2

@@ 2 Descent through Galois covers
splits. Thus there is a $K$-subgroup scheme $H$ of $E$ mapping isomorphically onto $[N]\phi(E)$. Then $G/H$ is an extension of $C/[N]\phi(E)$ by $G^{t}$, hence representable by a scheme $G^{\prime}$ with finitely generated component group. Letting $\overline{E}:=E/H$, we have $G/E\simeq G^{\prime}/\overline{E}$, and $\overline{E}(K_{s})$ has strictly smaller rank than $E(K_{s})$, hence $G^{\prime}/\overline{E}$ is almost-representable by induction. ∎
@@
with $H$ and $Q$ locally finite type group schemes over $K$. Then $\mathscr{G}$ is a smooth $K$-group scheme, and $\pi$ is a smooth morphism.
@@
We first show that $\mathscr{G}$ is a scheme. By [Stacks, Prop. 79.10.3], it suffices to show that $\mathscr{G}$ is a separated algebraic space over $K$. First, replacing $H$ and $Q$ by their maximal smooth $K$-subgroup schemes has no effect on the resulting sheaves, so we may assume that both groups are smooth. Choose an étale surjection $f\colon X\rightarrow Q$ such that $f$ lifts to a point $s\in\mathscr{G}(X)$. Then $Y:=\mathscr{G}\times_{Q}X\simeq H\times X$ via $(g,x)\mapsto(gs(x)^{-1},x)$, so this gives a (locally finite type) $K$-scheme admitting an étale surjection onto $\mathscr{G}$. It only remains to show that $\mathscr{G}$ is separated.
@@
is a closed embedding. Now $Y\times_{\mathscr{G}}Y$ is the subscheme of $Y\times Y=(H\times X)^{2}$ defined by the condition $h_{1}s(x_{1})=h_{2}s(x_{2})$. This condition in particular implies that $f(x_{1})=f(x_{2})$, so $Y\times_{\mathscr{G}}Y$ is contained in the preimage $Z\subset Y\times Y$ of the diagonal under the map $Y\times Y\rightarrow Q\times Q$. Because $Q$ – being a locally finite type group scheme over a field – is separated, $Z$ is a closed subscheme of $Y\times Y$. Finally, $Y\times_{\mathscr{G}}Y$ is the preimage of $1\in H(K)$ under the map $Z\rightarrow H$, $(h_{1},x_{1},h_{2},x_{2})\mapsto h_{1}s(x_{1})(h_{2}s(x_{2}))^{-1}$, hence it is a closed subscheme of $Z$.
@@
The map $q_{G}$ is surjective with kernel isomorphic to $E_{G}$, hence the same holds for $\zeta$. Lemma 2.10 therefore implies that $W$ is a smooth $F$-group scheme. Furthermore, the subsheaf $\eta^{-1}(E_{G})\subset W$ surjects onto $\ker(\phi)$ via $q_{F}\circ\zeta$. Call this surjection $\delta$. Then $\ker(\delta)=\zeta^{-1}(E_{F})$. This is étale with finitely generated component group, because $\ker(\zeta)\simeq\ker(q_{G})=E_{G}$. By Lemma 2.9, therefore, in order to prove the lemma it is enough to show that $U:=\eta^{-1}(E_{G})\subset W$ is a smooth $K$-group scheme with finitely generated component group.
@@
Proof.
Lemma 2.14.
A commutative extension of an almost-representable group sheaf by an affine, finite type group sheaf is almost-representable.
@@
By the effectivity of fpppf descent for relatively affine schemes, $W$ is a (smooth) $G$-scheme. Furthermore, the middle row shows that $W$ has finitely generated component group. It follows that $\mathscr{F}\simeq W/E$ is almost-representable. ∎
@@
We thus obtain an exact sequence of sheaves
@@ 3 Boundedness
(Boundedness) Let $Z$ be a $K$-scheme, and let $V\subset\Gamma(Z,\mathcal{O}_{Z})$ be a finite-dimensional $K$-vector space. Suppose given a $K$-scheme $T$ and $s\in\Gamma(Z\times_{K}T,\mathcal{O}_{Z\times_{K}T})$. We say that the pair $(T,s)$ is $V$-bounded when $s\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$.
@@
We first prove the representability. Once one has specified $g$, to give a $(\phi,V)$-bounded isomorphism $\psi$ is the same as giving a unit $u\in\Gamma(Z\times_{K}T,{\mathbf{G}}_{\rm{m}})$ such that $u,u^{-1}\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$. The datum of an element $f\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$ is the same as a $K$-morphism from $T$ to the vector group $\underline{V}$ defined by $V$. Thus to specify $\psi$ is the same as specifying a pair $(f,f^{\prime})$ of $K$-morphisms $T\rightarrow\underline{V}$ such that $ff^{\prime}=1$. We claim that this condition defines a closed subscheme $Y\subset\underline{V}\times\underline{V}$, in which case we see that ${\rm{Isom}}_{V}=S\times Y$ is a scheme.
@@
Suppose given an fpqc morphism of connected normal $K$-schemes $f\colon Y\rightarrow X$. Assume that ${\rm{DesCan}}_{Y/X}$ is almost-representable, and choose an isomorphism ${\rm{DesCan}}_{Y/X}\xrightarrow{\sim}G/E$ exhibiting it as such. Choose an étale cover $U\rightarrow G$ such that the element of ${\rm{DesCan}}_{Y/X}(G)$ corresponding to the quotient map $G\rightarrow G/E$ is represented on $U$ by a line bundle
@@
where $C:=\pi_{13}^{*}(\phi)\pi_{23}^{*}(\phi)^{-1}\pi_{12}^{*}(\phi^{-1})$. Choose a finite-dimensional subspace $H\subset\Gamma(Y\times_{X}Y\times_{X}Y,\mathcal{O})$ containing $C\cdot\pi_{13}^{*}(V)$ and $\pi_{12}^{*}(V)\cdot\pi_{23}^{*}(V)$, and choose a basis $\mathscr{B}$ for $H$. Then (3.2) reduces to the equations equating the coefficients of each element of $\mathscr{B}$ on both sides of (3.2). This condition defines a closed subscheme of ${\rm{Isom}}_{V}\times U$, as claimed.
We have thus obtained for each finite-dimensional $W\subset\Gamma(Y\times_{X}Y,\mathcal{O})$ a finite type $U$-scheme $r\colon S^{\prime}_{W}\rightarrow U$ which is a subfunctor of the functor of all descent data for $\mathscr{L}|Y\times_{K}U$, in such a manner that the functor of all descent data is the filtered direct limit of the $S^{\prime}_{W}$, and such that, for each $i$, there is a finite-dimensional $V_{i}$ such that, whenever $V_{i}\subset W$, $S^{\prime}_{W,i}:=S^{\prime}_{W}|r^{-1}(U_{i})$ defines all descent data on $\mathscr{L}|Y\times_{K}U_{i}$ over any smooth $K$-scheme $T$. Replacing $S^{\prime}_{W,i}$ by its maximal geometrically reduced closed subscheme $S_{W,i}$ (see [CGP, Lem.C.4.1]) does not alter its $T$-points for smooth $T$. Since $S^{\prime}_{W_{1},i}\subset S^{\prime}_{W_{2},i}$ is a closed subscheme with the same $K_{s}$-points when $V_{i}\subset W_{1}\subset W_{2}$, we see that $S_{W,i}$ is independent of $W\subset V_{i}$. We denote this common geometrically reduced finite type $U_{i}$-scheme by $S_{i}$. The $S_{i}$ together to yield a finite type $U$-scheme $S$ that is geometrically reduced over $K$ and which computes the functor of descent data for $\mathscr{L}$ from $Y$ to $X$ over any smooth $K$-scheme $T$.
@@
We claim that $H:={\rm{Units}}_{Y/K}/{\rm{Units}}_{X/K}$ is a smooth $K$-group scheme. Indeed, to verify this, it suffices to prove that the map on component groups induced by the inclusion
@@
is injective. For this, we may pass to a finite Galois extension and thereby suppose that we have a unit $u$ on $X$ that does not live in ${\rm{Units}}_{X/K}^{0}$, and we must show that its pullback to $Y$ does not live in ${\rm{Units}}_{Y/K}^{0}$. By Proposition 2.1, $u$ is in the identity component of $X$ precisely when it is algebraic over $K$, and similarly for units on $Y$. Thus what we must show is that, if $u$ is not algebraic, then neither is its pullback to $Y$. But this follows from the fact that the pullback map $f^{*}\colon\Gamma(X,\mathcal{O}_{X})\rightarrow\Gamma(Y,\mathcal{O}_{Y})$ is injective, because $f$ is faithfully flat.
It follows that $\mathscr{F}$ is computed by the geometrically reduced finite type $U$-algebraic space $\mathscr{S}:=S/H$ with structure map $r^{\prime}\colon\mathscr{S}\rightarrow U$. Furthermore, $\mathscr{F}_{i}$ is computed by the finite type $U_{i}$-algebraic space $\mathscr{S}^{\prime}_{W,i}:=S^{\prime}_{W,i}/H$ for any finite-dimensional $V\subset W_{i}$. Let $R\subset U\times U$ be the étale equivalence relation such that $U/R=G$. Because $G$ is a locally finite type $K$-group scheme, it is separated, so $R\subset U\times U$ is a closed subscheme. Then $\iota\colon R^{\prime}:=(r^{\prime}\times r^{\prime})^{-1}(R)\hookrightarrow\mathscr{S}\times\mathscr{S}$ is an et́ale equivalence relation and a closed embedding, and we let $\mathscr{X}:=\mathscr{S}/R^{\prime}$, a locally finite type, separated (because $\iota$ is a closed embedding), geometrically reduced $K$-algebraic space that computes the sheaf $\mathscr{G}$ on the smooth-étale site of ${\rm{Spec}}(K)$ which sits in the following Cartesian diagram
@@
Let $D_{i}:=a(U_{i})\subset{\rm{DesCan}}_{Y/X}$, an open subset. The group structure on $U/R=G$, together with that on (isomorphism classes of) line bundles on $X$ descending $\mathscr{L}$, yields a multiplication map for any $i,j,k$ and $W_{1},W_{2}$
@@
where ${\bf k}\subset I$ is a finite subset such that the union of the $U_{k}$ contains $D_{i}\cdot D_{j}$, and $W_{3}$ is sufficiently large depending on $W_{1},W_{2}$ (and $\mathscr{X}_{W_{3},{\bf k}}$ has the obvious meaning). In this manner one obtains a group structure on the colimit (over all $W,i$) of the $\mathscr{X}^{\prime}_{W,i}$, hence on the colimit of the maximal geometrically reduced closed subschemes, which is $\mathscr{X}$. Thus $\mathscr{X}$ is a geometrically reduced, separated, locally finite type $K$-group algebraic space, so it is a smooth commutative $K$-group scheme [Stacks, Prop. 79.10.3]. Because the map $m\colon\mathscr{X}\rightarrow G$ is of finite type, $\mathscr{X}$ has finitely generated component group. Thus, $\operatorname{Pic}^{+}_{X/K}\simeq\mathscr{X}/E$ is almost-representable. ∎
@@
Working locally, we may assume that $T={\rm{Spec}}(R)$ and that $u$ lies in a single component $C$ of ${\rm{Units}}_{Y/K}$. Choose a unit $v$ on $Y$ defining a point in $C(K)$. Then $C(R)=u(R\otimes_{K}B)^{\times}$, so $v$ lies in the $R$-module $u(R\otimes_{K}B)$. By descent theory, an element of $u(R\otimes_{K}B)$ is the pullback of a section of $X_{R}$ precisely when its pullbacks along the two projections $(Y\times_{X}Y)_{R}\rightarrow Y_{R}$ agree. Because $R$ is $K$-flat, the elements of $u(R\otimes_{K}B)$ whose pullbacks agree are $R\otimes_{K}M$, where $M\subset uB$ are the elements of $uB$ whose pullbacks agree – that is, that descend to sections of $\Gamma(X,\mathcal{O}_{X})$.
If $M=0$, then $v=0$, a contradiction, so $M\neq 0$. Let $0\neq m\in M$. Then $f^{*}(m)\in uB\backslash\{0\}$. Because $Y$ is connected and reduced, $B$ is a field, so $f^{*}(m)$ is a unit of $Y$, hence $m$ is a unit of $X$. Replacing $u$ by $f^{*}(m)$, therefore, we may assume that $u$ descends to a unit of $X$, hence lives in ${\rm{Units}}_{X/K}$. So modifying by $u$, we have reduced to the case in which $C={\rm{Units}}_{Y/K}^{0}=\operatorname{R}_{B/K}({\mathbf{G}}_{\rm{m}})$. Then $M=A$, so an element of $(R\otimes_{K}B)^{\times}$ which is pulled back from $X$ lies in $(R\otimes_{K}A)^{\times}={\rm{Units}}_{X/K}^{0}(R)$. ∎
4 Descent through purely inseparable extensions
In this section we will prove the results required in order to descend the almost-representability of the restricted Picard functor through purely inseparable extensions (although we postpone the proof that representability descends through such extensions until the next section). This may be regarded as the technical heart of the entire paper. We will proceed by verifying the hypotheses of Proposition 3.4. We begin by verifying the almost-representability of ${\rm{DesCan}}$ in this setting.

### v4

@@ 2 Descent through Galois covers
splits. Thus there is a $K$-subgroup scheme $F^{\prime}$ of $F$ mapping isomorphically onto $[N]\phi(F)$. Then $H/F^{\prime}$ is an extension of $C/[N]\phi(F)$ by $H^{t}$, hence ([Stacks, Tag 0B8G] again) representable by a scheme $G^{\prime}$ with finitely generated component group. Letting $\overline{F}:=F/F^{\prime}$, we have $H/F\simeq G^{\prime}/\overline{F}$, and $\overline{F}(K_{s})$ has strictly smaller rank than $F(K_{s})$, hence $G^{\prime}/\overline{F}$ has the form asserted in the lemma by induction. ∎
@@
with $H$ and $Q$ locally finite type group schemes over $K$. Then $\mathscr{G}$ is represented by a smooth $K$-group scheme, and $\pi$ is a smooth morphism.
@@
We first show that $\mathscr{G}$ is a scheme. By [Stacks, Tag 0B8G], it suffices to show that $\mathscr{G}$ is a separated algebraic space over $K$. First, replacing $H$ and $Q$ by their maximal smooth $K$-subgroup schemes has no effect on the resulting sheaves, so we may assume that both groups are smooth. Choose an étale surjection $f\colon X\rightarrow Q$ such that $f$ lifts to a point $s\in\mathscr{G}(X)$. Then $Y:=\mathscr{G}\times_{Q}X\simeq H\times X$ via $(g,x)\mapsto(gs(x)^{-1},x)$, so this gives a (locally finite type) $K$-scheme admitting an étale surjection onto $\mathscr{G}$. It remains to prove that $\mathscr{G}$ is separated.
@@
is a closed embedding. Now $Y\times_{\mathscr{G}}Y$ is the subscheme of $Y\times Y=(H\times X)^{2}$ defined by the condition $h_{1}s(x_{1})=h_{2}s(x_{2})$. This condition in particular implies that $f(x_{1})=f(x_{2})$, so $Y\times_{\mathscr{G}}Y$ is contained in the preimage $Z\subset Y\times Y$ of the diagonal under the map $Y\times Y\rightarrow Q$. Because $Q$ – being a locally finite type group scheme over a field – is separated, $Z$ is a closed subscheme of $Y\times Y$. Finally, $Y\times_{\mathscr{G}}Y$ is the preimage of $1\in H(K)$ under the map $Z\rightarrow H$, $(h_{1},x_{1},h_{2},x_{2})\mapsto h_{1}s(x_{1})(h_{2}s(x_{2}))^{-1}$, hence it is a closed subscheme of $Z$.
@@
The map $q_{G}$ is surjective with kernel isomorphic to $E_{G}$, hence the same holds for $\zeta$. Lemma 2.10 therefore implies that $W$ is a smooth $F$-group scheme. Furthermore, the subsheaf $\eta^{-1}(E_{G})\subset W$ surjects onto $\ker(\phi)$ via $q_{F}\circ\zeta$. Call this surjection $\delta$. Then $\ker(\delta)=\zeta^{-1}(E_{F})$. This is an extension of $E_{F}$ by $E_{G}$, hence étale with finitely generated component group. In order to prove the lemma, therefore, it is enough to show that $U:=\eta^{-1}(E_{G})\subset W$ is a smooth $K$-group scheme with finitely generated component group.
@@
A commutative extension of an almost-representable group sheaf by a finite type $K$-group scheme with finitely generated component group is almost-representable.
@@
By Lemma 2.10, $W$ is a (smooth) $G$-scheme. Furthermore, the middle row shows that $W$ has finitely generated component group. It follows that $\mathscr{F}\simeq W/E$ is almost-representable. ∎
@@
By sheafifying the low-degree exact sequence associated to this spectral sequence, we obtain an exact sequence of sheaves
@@ 3 Boundedness
(Boundedness) Let $Z$ be a qcqs $K$-scheme, and let $V\subset\Gamma(Z,\mathcal{O}_{Z})$ be a finite-dimensional $K$-vector space. Suppose given a $K$-scheme $T$ and $s\in\Gamma(Z\times_{K}T,\mathcal{O}_{Z\times_{K}T})$. We say that the pair $(T,s)$ is $V$-bounded when $s\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$.
@@
We first prove the representability. When $Z=\emptyset$, the assertion is clear, so assume $Z\neq\emptyset$. Once one has specified $g$, to give a $(\phi,V)$-bounded isomorphism $\psi$ is the same as giving a unit $u\in\Gamma(Z\times_{K}T,{\mathbf{G}}_{\rm{m}})$ such that $u,u^{-1}\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$. The datum of an element $f\in V\otimes_{K}\Gamma(T,\mathcal{O}_{T})$ is the same as a $K$-morphism from $T$ to the vector group $\underline{V}$ defined by $V$. Thus to specify $\psi$ is the same as specifying a pair $(f,f^{\prime})$ of $K$-morphisms $T\rightarrow\underline{V}$ such that $ff^{\prime}=1$. We claim that this condition defines a closed subscheme $Y\subset\underline{V}\times\underline{V}$, in which case we see that ${\rm{Isom}}_{V}=S\times Y$ is a scheme.
@@
Suppose given an fpqc morphism of connected normal finite type $K$-schemes $f\colon Y\rightarrow X$. Assume that ${\rm{DesCan}}_{Y/X}$ is almost-representable, and choose an isomorphism ${\rm{DesCan}}_{Y/X}\xrightarrow{\sim}G/E$ exhibiting it as such. Choose an étale cover $U\rightarrow G$ such that the element of ${\rm{DesCan}}_{Y/X}(G)$ corresponding to the quotient map $G\rightarrow G/E$ is represented on $U$ by a line bundle
@@
where $C:=\pi_{13}^{*}(\phi)\pi_{23}^{*}(\phi^{-1})\pi_{12}^{*}(\phi^{-1})$. Choose a finite-dimensional subspace $H\subset\Gamma(Y\times_{X}Y\times_{X}Y,\mathcal{O})$ containing $C\cdot\pi_{13}^{*}(V)$ and $\pi_{12}^{*}(V)\cdot\pi_{23}^{*}(V)$, and choose a basis $\mathscr{B}$ for $H$. Then (3.2) reduces to the equations equating the coefficients of each element of $\mathscr{B}$ on both sides of (3.2). This condition defines a closed subscheme of ${\rm{Isom}}_{V}\times U$, as claimed.
We have thus obtained for each finite-dimensional $W\subset\Gamma(Y\times_{X}Y,\mathcal{O})$ a finite type $U$-scheme $r\colon S^{\prime}_{W}\rightarrow U$ which is a subfunctor of the functor of all descent data for $\mathscr{L}|Y\times_{K}U$, in such a manner that the functor of all descent data is the filtered direct limit of the $S^{\prime}_{W}$, and such that, for each $i$, there is a finite-dimensional $V_{i}$ such that, whenever $V_{i}\subset W$, $S^{\prime}_{W,i}:=S^{\prime}_{W}|r^{-1}(U_{i})$ defines all descent data on $\mathscr{L}|Y\times_{K}U_{i}$ over any smooth $K$-scheme $T$. Replacing $S^{\prime}_{W,i}$ by its maximal geometrically reduced closed subscheme $S_{W,i}$ (see [CGP, Lem.C.4.1]) does not alter its $T$-points for smooth $T$. Since $S^{\prime}_{W_{1},i}\subset S^{\prime}_{W_{2},i}$ is a closed subscheme with the same $K_{s}$-points when $V_{i}\subset W_{1}\subset W_{2}$, we see that $S_{W,i}$ is independent of $W\subset V_{i}$. We denote this common geometrically reduced finite type $U_{i}$-scheme by $S_{i}$. The $S_{i}$ glue together to yield a finite type $U$-scheme $S$ that is geometrically reduced over $K$ and which computes the functor of descent data for $\mathscr{L}$ from $Y$ to $X$ over any smooth $K$-scheme $T$ with a map $T\rightarrow U_{i}$.
@@
We claim that $H:={\rm{Units}}_{Y/K}/{\rm{Units}}_{X/K}$ is a smooth $K$-group scheme. To verify this, we claim that it suffices to prove that the map on component groups induced by the inclusion
@@
is injective. Indeed, letting $E_{X},E_{Y}$ denote the components groups of ${\rm{Units}}_{X/K}$ and ${\rm{Units}}_{Y/K}$, respectively, this injectivity would yield by the snake lemma an exact sequence
$0\longrightarrow{\rm{Units}}_{Y/K}^{0}/{\rm{Units}}_{X/K}^{0}\longrightarrow{\rm{Units}}_{Y/K}/{\rm{Units}}_{X/K}\longrightarrow E_{Y}/E_{X}\longrightarrow 0,$
and representability by a scheme then follows as usual from [Stacks, Tag 0B8G].
To prove the claimed injectivity on component groups, we may pass to a finite Galois extension and thereby suppose that we have a unit $u$ on $X$ that does not live in ${\rm{Units}}_{X/K}^{0}$, and we must show that its pullback to $Y$ does not live in ${\rm{Units}}_{Y/K}^{0}$. By Proposition 2.1, $u$ is in the identity component of $X$ precisely when it is algebraic over $K$, and similarly for units on $Y$. Thus what we must show is that, if $u$ is not algebraic, then neither is its pullback to $Y$. But this follows from the fact that the pullback map $f^{*}\colon\Gamma(X,\mathcal{O}_{X})\rightarrow\Gamma(Y,\mathcal{O}_{Y})$ is injective, because $f$ is faithfully flat.
It follows that $\mathscr{F}$ is computed by the geometrically reduced finite type $U$-algebraic space $\mathscr{S}:=S/H$ with structure map $r^{\prime}\colon\mathscr{S}\rightarrow U$. Furthermore, $\mathscr{F}_{i}:=h^{-1}(U_{i})$ is computed by the finite type $U_{i}$-algebraic space $\mathscr{S}^{\prime}_{W,i}:=S^{\prime}_{W,i}/H$ for any finite-dimensional $V\subset W_{i}$. Let $R\subset U\times U$ be the étale equivalence relation such that $U/R=G$. Because $G$ is a locally finite type $K$-group scheme, it is separated, so $R\subset U\times U$ is a closed subscheme. Then $\iota\colon R^{\prime}:=(r^{\prime}\times r^{\prime})^{-1}(R)\hookrightarrow\mathscr{S}\times\mathscr{S}$ is an et́ale equivalence relation and a closed embedding, and we let $\mathscr{X}:=\mathscr{S}/R^{\prime}$, a locally finite type, separated (because $\iota$ is a closed embedding), geometrically reduced $K$-algebraic space that computes the sheaf $\mathscr{G}$ on the smooth-étale site of ${\rm{Spec}}(K)$ which sits in the following Cartesian diagram
@@
Let $D_{i}:=a(U_{i})\subset{\rm{DesCan}}_{Y/X}$. The group structure on $U/R=G$, together with that on (isomorphism classes of) line bundles on $X$ descending $\mathscr{L}$, yields a multiplication map for any $i,j,k$ and $W_{1},W_{2}$
@@
where ${\bf k}\subset I$ is a finite subset such that the union of the $D_{k}$ contains $D_{i}\cdot D_{j}$, and $W_{3}$ is sufficiently large depending on $W_{1},W_{2}$ (and $\mathscr{X}_{W_{3},{\bf k}}$ has the obvious meaning). In this manner one obtains a group structure on the colimit (over all $W,i$) of the $\mathscr{X}^{\prime}_{W,i}$, hence on the colimit of the maximal geometrically reduced closed subspaces, which is $\mathscr{X}$. Thus $\mathscr{X}$ is a geometrically reduced, separated, locally finite type $K$-group algebraic space, so it is a smooth commutative $K$-group scheme [Stacks, Tag 0B8G]. Because the map $m\colon\mathscr{X}\rightarrow G$ is of finite type, $\mathscr{X}$ has finitely generated component group. Thus, $\operatorname{Pic}^{+}_{X/K}\simeq\mathscr{X}/E$ is almost-representable, where the inclusion of $E$ in $\mathscr{X}$ is via the inclusion $E\hookrightarrow\mathscr{G}$, and the fact that the restriction of $\mathscr{X}$ to the smooth-étale site of ${\rm{Spec}}(K)$ (of which $E$ is an object) computes $\mathscr{G}$. ∎
@@
Working locally, we may assume that $T={\rm{Spec}}(R)$ and that $u$ lies in a single component $C$ of ${\rm{Units}}_{Y/K}$. Choose a unit $v$ on $Y$ defining a point in $C(K)$. Then $C(R)=v(R\otimes_{K}B)^{\times}$, so $u$ lies in the $R$-module $v(R\otimes_{K}B)$. By descent theory, an element of $v(R\otimes_{K}B)$ is the pullback of a section of $X_{R}$ precisely when its pullbacks along the two projections $(Y\times_{X}Y)_{R}\rightarrow Y_{R}$ agree. Because $R$ is $K$-flat, the elements of $v(R\otimes_{K}B)$ whose pullbacks agree are $R\otimes_{K}M$, where $M\subset vB$ are the elements of $vB$ whose pullbacks agree – that is, that descend to sections of $\Gamma(X,\mathcal{O}_{X})$.
If $M=0$, then $u=0$, a contradiction, so $M\neq 0$. Let $0\neq m\in M$. Because $Y$ is connected and reduced, $B$ is a field, so $m$ is a unit of $Y$, hence $m$ is a unit of $X$. Replacing $v$ by $m$, therefore, we may assume that $v$ descends to a unit of $X$, hence lives in ${\rm{Units}}_{X/K}$. So modifying by $v$, we have reduced to the case in which $C={\rm{Units}}_{Y/K}^{0}=\operatorname{R}_{B/K}({\mathbf{G}}_{\rm{m}})$. Then $M=A$, so an element of $(R\otimes_{K}B)^{\times}$ which is pulled back from $X$ lies in $(R\otimes_{K}A)^{\times}={\rm{Units}}_{X/K}^{0}(R)$. ∎
4 Descent through radicial covers
In this section we will prove the results required in order to descend the almost-representability of the restricted Picard functor through radicial covers (although we postpone the proof that representability descends through such covers until the next section). This may be regarded as the technical heart of the entire paper. We will proceed by verifying the hypotheses of Proposition 3.4. We begin by identifying ${\rm{DesCan}}$ in this setting.